OpenStudy (manray):

Anderson uses the discriminant to correctly find the number of real solutions of the quadratic equation 1/2x2 + 4x + 8 = 0. Which explanation could Anderson provide?

likeabossssssss (likeabossssssss):

do u got a answer?

OpenStudy (manray):

The equation has no real number solutions because the discriminant is 0. The equation has one real number solution because the discriminant is 0. The equation has no real number solutions because the discriminant is less than 0. The equation has two real number solutions because the discriminant is greater than 0.

alones (alones):

Any ideas?

likeabossssssss (likeabossssssss):

^^

OpenStudy (manray):

i was thinking A

OpenStudy (sprattfield):

The answer is B

likeabossssssss (likeabossssssss):

no direct answer plz

likeabossssssss (likeabossssssss):

@sprattfield

alones (alones):

Since @sprattfield Knows the answer, I'd like them to answer it (;

OpenStudy (sprattfield):

Here is your explanation: The discriminant equation is b^2-4ac b=4 a=1 c=8 4^2-4(1)(8) = 16 - 32 = -16 -16 < 0, thus the answer is B

OpenStudy (manray):

thanks guys!!!

OpenStudy (sprattfield):

No problem

likeabossssssss (likeabossssssss):

oh yea welcome to os @ManRay btw may u close this

likeabossssssss (likeabossssssss):

thank u :)